four-sevenths of a school's faculty are women. four-fifths of the male faculty members are married, and nine of the male faculty are unmarried. how many faculty members are there?
Answer by IWork4Dessert(57)

Let's work this from the end to the start. You know two things about the male faculty members, so begin with those. 4/5 of the men are married, and 9 of them are not. This means that 1/5 of the men are unmarried, so 9 is 1/5 of the total amount of men. Now you can set up a proportion:
1/5=9/x
Solve the proportion by cross-multiplying.
x=45
There are 45 men total. You know now that since 4/7 of the faculty are women, then 3/7 of them are men. So you have another proportion:
3/7=45/x
Cross-multiply again.
3x=315
Divide both sides by 3 to get the variable(letter) alone.
x=105
105 people are in the faculty total.